OpenStudy (mateeldaa):

I've been stuck on this problem for about 30 minutes now. Any help is greatly appreciated!! Solve the equation. (2/(z-5))-((6/4z+9))+((3/z-5))=0

OpenStudy (perl):

is that (i am going to remove some parentheses) 2/(z-5)-6/(4z+9)+3/(z-5)=0

OpenStudy (mateeldaa):

\[\frac{ 2 }{ z-5 }-\frac{ 6 }{ 4z+9 }+\frac{ 3 }{ z-5 } = 0\]

OpenStudy (perl):

so before we do anything, notice that the left side expression is not defined if z = 5, or z = -9/4 , because of division by zero

OpenStudy (mateeldaa):

right, since dividing by zero gives you an undefined

OpenStudy (perl):

there are two ways to solve this, the easy way is to multiply both sides by the common denominator (z-5)*(4z + 9) , that will cancel the denominators

OpenStudy (mateeldaa):

I tried it that way, but I multiplied it by the one on the left first and then the one on the right and that's where I got stuck

OpenStudy (perl):

Multiply both sides by [(z-5) * (4z+9 )] [(z-5) * (4z+9 )] * [2/(z-5)-6/(4z+9)+3/(z-5) ] =0*[(z-5) * (4z+9 )] Did you get 2*(4z+9) - 6 ( z-5) + 3(4z+9) = 0

OpenStudy (perl):

if not, i can explain how i canceled

OpenStudy (mateeldaa):

No, I don't see how you got that

OpenStudy (perl):

ok we need to distribute each term

OpenStudy (perl):

the right side is clearly zero, since zero times anything is zero , agreed?

OpenStudy (mateeldaa):

agreed

OpenStudy (perl):

[(z-5) * (4z+9 )] * [2/(z-5)-6/(4z+9)+3/(z-5) ] = 0 distribute [(z-5)*(4z+9)] * 2/(z-5) + [(z-5)*(4z+9)] *(-6)/(4z+9) + [(z-5)*(4z+9)]*3/(z-5) = 0

OpenStudy (perl):

the general form of this is [u ] * ( a - b + c ) = u*a + u*(-b) + u*c

OpenStudy (mateeldaa):

so when you distribute the denominator it'll cancel out and you'll be left with distributing the top number?

OpenStudy (perl):

right

OpenStudy (perl):

for example when you multiply [(z-5)*(4z+9)] * 2/(z-5) the (z-5)/(z-5) cancels,

OpenStudy (perl):

so after canceling, you should get 2*(4z+9) - 6 ( z-5) + 3(4z+9) = 0

OpenStudy (mateeldaa):

Okay, I got there

OpenStudy (perl):

ok now distribute

OpenStudy (perl):

2*(4z+9) - 6 ( z-5) + 3(4z+9) = 0 distribute: 2*4z + 2*9 - 6*z -6*(-5) + 3*4z+3*9 = 0

OpenStudy (mateeldaa):

8z+18-6z+30+12z+27=0

OpenStudy (perl):

looks good

OpenStudy (perl):

now collect like terms

OpenStudy (mateeldaa):

14z+75=0

OpenStudy (perl):

now solve for z

OpenStudy (mateeldaa):

z=-75/14

OpenStudy (perl):

right

OpenStudy (perl):

now test this value, plug it back into the original equation

OpenStudy (perl):

actually you can skip this part, (i just checked and its correct)

OpenStudy (perl):

but you could if you wanted to :)

OpenStudy (mateeldaa):

Thanks so much!!!